x/2 years ago,Jason was 12,and x/2years from now he will be 2x years old,how old will he be 3x years from now
Answer by richwmiller(17219)

j-x/2=12;
j+x/2=2x;
k=j+3x
j=18
x=12
k=54
check
18-12/2=12
ok
18+6=2*12
24=24
ok
18+3*12=18+36=54
ok